# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ C.1. Proof of Proposition 6 [No Type (2) edge implies non-$3$-SIP map]
By Lemma 25, every graph-nonedge pair in Figure 12 has a proper non-$3$-SIP map. It remains to show that these are the only minimal graph-nonedge pairs with no Type (2) edge. Consider any graph-nonedge pair $(G,f)$, where $f=uv$. If $(G,f)$ is a pair in Figure 12, then it is easy to verify that $(G,f)$ is minimal and has no Type (2) edge. Next, we show that if $(G,f)$ is minimal and has no Type (2) edge, then it is some pair in Figure 12. Since $(G,f)$ is minimal, $G\cup f$ is an atom and has an $f$-preserving forbidden minor $[G\cup f]$. Furthermore, Statements (i)-(iv) of Lemma 23 are true and show that $5\leq|V(G)|\leq 8$. Let $H=[G\cup f]\setminus\{[u],[v]\}$ and consider the subgraph $H^{-1}$ of $G\cup f$. If $|[u]^{-1}|=|[v]^{-1}|=1$, then Statement (ii) shows that $(G,f)$ is the pair in either Figure 12(a) or 12(b). Hence, assume wlog that $|[u]^{-1}|>1$. By Statement (iv), we have $[u]^{-1}=\{u,x\}$. There are three cases.
@@
Let $w$ be the vertex in $H^{-1}$ that neighbors $x$ but not $u$ if such a vertex exists, or any vertex in $H^{-1}$ that neighbors $x$ otherwise. Also, let $J$ be the subgraph of $G\cup f$ induced by $[u]^{-1}$. By Statement (iii), $J\setminus x$ is connected and contains $u$. Hence, $xw$ is contracted in the minor obtained via the vertex exchange from $[u]^{-1}$ to $[w]^{-1}$ that exchanges $x$ and fixes $J\setminus x$, and this minor either is or has an $f$-preserving forbidden minor. However, this contradicts Statement (i).
@@
Next, consider the case where $|[v]^{-1}|=1$. If $xw$ is a nonedge of $G\cup f$, then $(G,f)$ is the pair in either Figure 12(c) or 12(d). Otherwise, $\{x,v,w\}$ is a clique separator of $G\cup f$, contradicting the fact that $G\cup f$ is an atom. Lastly, consider the case where $|[v]^{-1}|>1$. The discussion above shows that $w$ is the only neighbor of $v$ in $H^{-1}$. Let $w^{\prime}$ be any vertex in $H^{-1}\setminus w$ that neighbors $x$. Also, let $J$ be defined as in Case 2. Then, $xw^{\prime}$ is contracted in the minor obtained via the vertex exchange from $[u]^{-1}$ to $[w^{\prime}]^{-1}$ that exchanges $x$ and fixes $J\setminus x$, and this minor has an $f$-preserving forbidden minor. However, this contradicts Statement (i).
@@
C.2. Proof of Proposition 7 [Graphs with edge-separating minors have an expanded edge-separating cms]
The proof of Proposition 7 requires Lemma 28, below. This lemma is proved using Lemma 27, below, which in turn is proved using Lemma 26, also below.
@@ C.2. Proof of Proposition 7 [Graphs with edge-separating minors have an expanded edge-separating cms]
If an edge $e$ of a graph $G$ is preserved in some $K_{5}$ minor, then it is clearly preserved in some induced $K_{5}$ minor. Otherwise, if $e$ is preserved in some $K_{2,2,2}$ minor, then consider the minor $[G]$ obtained by performing all contractions between $G$ and this $K_{2,2,2}$ minor. If $[G]$ is $K_{2,2,2}$, then we are done. Otherwise, $[G]$ has $K_{2,2,2}$ as a proper spanning subgraph, in which case $[G]$ has an $[e]$-preserving $K_{5}$ minor. This completes the proof. ∎
Lemma 27 (Edge-preserving forbidden minors extend minors of atoms).
Proof.
Observe that $K_{5}$ and $K_{2,2,2}$ are both atoms. Hence, the lemma follows by applying Lemma 26 followed by Lemma 58 in Appendix E. ∎
Proof.
By definition, any $e$-separating forbidden minor $[G]$ has an $[e]$-preserving forbidden minor but none of its atoms that contain $[e]$ have such a minor. Hence, Lemma 27 shows that some atom $M$ of $[G]$ does not contain some endpoint $[u]$ of $[e]$ and has an $[e]$-preserving forbidden minor. The above facts imply the existence of some clique minimal $[ux]$-separator $C$ of $[G]$, for any vertex $[x]$ of $M\setminus C$, such that $M\setminus C$ is non-empty. Thus, $(C,M)$ is an $e$-separating pair of $[G]$, and the proof is complete. ∎
Proof of Proposition 7.
If $[G]$ is an $e$-separating minor, then Lemma 28 shows that it has an $e$-separating pair. Thus, $G$ clearly has a top-level expanded $e$-separating pair, and hence a top-level expanded $e$-separating cms. ∎
C.3. Proof outline for Proposition 8 [Minimal pair inductive structure]
The proof of Proposition 8 is very technical, and so we split it into the following four sections.
C.3.1. Proof of Lemma 5[Top-Level Structure]
Here we prove Lemma 5 from Section 4 using Lemmas 30 and 32, below. Lemma 30 is proved using Lemma 29, also below.
Let $C^{-1}$ be any expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, with $f=uv$, and let its expanded $f$-component and expanded minor component be $H^{-1}$ and $I^{-1}$, respectively. Also, consider the minor $[G\cup f]$ obtained by contracting any edge of $(H^{-1}\setminus f)\setminus C^{-1}$ and any $f$-preserving forbidden minor $[I^{-1}]_{A}$. There exists a pair $xy$ of distinct vertices in $C^{-1}$ that is preserved in $[I^{-1}]_{A}$ and some paths $P_{ux}$ and $P_{vy}$ in $G\cup f$ between $u$ and $x$ and between $v$ and $y$, respectively, that each contain exactly one vertex in $I^{-1}$ and such that $[P_{ux}]$ and $[P_{vy}]$ are vertex-disjoint.
Proof.
Since $f$ is preserved in $[I^{-1}]_{A}$ and $(G\cup f)\setminus C^{-1}$ is disconnected, there exists a pair of distinct vertices $x^{\prime}y^{\prime}$ in $C^{-1}$ and an extension $[G\cup f]_{B}$ of $[I^{-1}]_{A}$ such that $[f]_{B}=[x^{\prime}y^{\prime}]_{B}$. This implies that $x^{\prime}y^{\prime}$ is preserved in $[I^{-1}]_{A}$ and some vertex-disjoint paths $P_{ux^{\prime}}$ and $P_{vy^{\prime}}$ in $G\cup f$ between $u$ and $x^{\prime}$ and between $v$ and $y^{\prime}$, respectively, each contain exactly one vertex in $I^{-1}$. If $[P_{ux^{\prime}}]$ and $[P_{vy^{\prime}}]$ are vertex-disjoint, then setting $xy=x^{\prime}y^{\prime}$ proves the lemma.
@@ C.3.1. Proof of Lemma 5[Top-Level Structure]
Lemma 30 (No Type (1) edge in $(H^{-1}\setminus f)\setminus C^{-1}$).
Let $C^{-1}$ be any expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, with $f=uv$, and let its expanded $f$-component and expanded minor component be $H^{-1}$ and $I^{-1}$, respectively. For any minor $[G\cup f]$ obtained by contracting any edge of $(H^{-1}\setminus f)\setminus C^{-1}$, $[I^{-1}]$, and hence $[G\cup f]$, has an $[f]$-preserving forbidden minor.
Proof.
By definition, $I^{-1}$ has an $f$-preserving forbidden minor $[I^{-1}]_{A}$. Hence, since $I^{-1}$ is preserved in $[G\cup f]$, it suffices to show that the edge contracted in $[G\cup f]$ is contracted in some extension of $[I^{-1}]_{A}$. This follows easily from Lemma 29. ∎
Next, Lemma 32 is proved using Lemma 31, below, which is in turn proved using Lemma 4 from Section 4. We prove Lemma 4 below. An induced subgraph $H$ of a graph $G$ is retained in a minor $[G]$ if each pair of vertices in $H$ is retained in $[G]$.
@@
For Statement (1), since $(G\cup f)\setminus C^{-1}$ is disconnected and $G\cup f$ is an atom, $C^{-1}$ is not a clique. Hence, Figure 13 shows all cases for $C^{-1}$ if $|V(C^{-1})|\leq 5$. We complete the proof by showing that $|V(C^{-1})|\leq 5$. Assume to the contrary that this is not the case. Consider any $f$-separating minor $[G\cup f]_{A}$ such that $(C,M)$ is one of its $f$-separating pairs. Then, $C$ is a clique on at least five vertices. Consequently, since any atom $J$ of $[G\cup f]_{A}$ that contains $[f]_{A}$ has no $[f]_{A}$-preserving forbidden minor, by definition, $J$ cannot contain $C$. This implies that there exists a clique minimal $[ux]_{A}$-separator $E$ of $[G\cup f]_{A}$, for some endpoint $[u]_{A}$ of $[f]_{A}$ and any vertex $[x]_{A}$ in $C\setminus E$, such that $C\setminus E$ is non-empty. Clearly, $(E^{-1},M^{-1})$ is an expanded $f$-separating pair of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$. However, this shows that $C^{-1}$ is not in the top-level of $(G,f)$, which is a contradiction. Therefore, we have $|V(C^{-1})|\leq 5$, and so Statement (1) is proved.
Next, for Statement (2), observe that $C^{-1}$ is retained in $[H^{-1}\cup I^{-1}]$, and $[H^{-1}]$ and $[I^{-1}]$ are vertices of $[H^{-1}\cup I^{-1}]$. Hence, Statement (1) shows that $[C^{-1}]$ is one the graphs in Figure 13. Consequently, $[H^{-1}\cup I^{-1}]$ is some graph $G^{\prime}$ that can be obtained from one of the graphs in Figure 14 by deleting some edges, each of which has one endpoint in $\{a,b\}$ and endpoint in $\{w_{i}\}$. Furthermore, $[C^{-1}]$ is the subgraph of $G^{\prime}$ induced by the set $\{w_{i}\}$. If $G^{\prime}$ can be obtained without deleting any edges, then we are done. Otherwise, assume wlog that the above-mentioned minor $[G\cup f]_{A}$ is obtained by contracting $w_{1}w_{2}$. It is easy to see that either $G\cup f$ is not an atom or, for any endpoint $[u]_{A}$ of $[f]_{A}$ and any vertex $[x]_{A}$ of $M\setminus C$, $C$ is not a minimal $[ux]_{A}$-separator of $[G\cup f]_{A}$. In either case, we get a contradiction, and thus Statement (2) is proved. ∎
@@
Lemma 32 (No Type (2) edge in $(H^{-1}\setminus f)\setminus C^{-1}$).
Proof.
@@
Finally, we prove Lemma 5.
Proof of Lemma 5.
Since $w$ is a vertex of $H^{-1}\setminus(C^{-1}\cup\{u,v\})$, it must be adjacent to some vertex $z$ in $H^{-1}\setminus C^{-1}$. Note that Lemmas 30 and 32 imply that $wz$ is of Type (3). Hence, $z$ must be either $u$ or $v$, and both $wu$ and $wv$ are edges. Next, since $G\cup f$ is an atom, $w$ must be adjacent to some vertex other than $u$ or $v$. The same two lemmas imply that this vertex is contained in $C^{-1}$. Thus, the lemma is proved. ∎
C.3.2. Proof of Lemma 6 [Pruned possible graphs for $C^{-1}$]
Here we prove Lemma 6 using Lemma 34, below, which is in turn proved using Lemma 33, also below.
@@ C.3.2. Proof of Lemma 6 [Pruned possible graphs for $C^{-1}$]
Figure 19. See Lemma 34. All possible expanded $f$-components of an expanded $f$-separating cms that has exactly five vertices and is in the top-level of a minimal graph-nonedge pair $(G,f)$. Dashed line-segments may be edges or nonedges.
Lemma 34 (Cases for $H^{-1}$ when $C^{-1}$ has $5$ vertices).
Let $C^{-1}$ be a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$. If $|V(C^{-1})|=5$, then the expanded $f$-component of $C^{-1}$ is one of the graphs in Figure 19 such that $f$ is the red edge and $C^{-1}$ is the subgraph induced by the set $\{w_{i}\}$.
Proof.
Consider any $f$-separating minor $[G\cup f]$ obtained by contracting a single edge of $C^{-1}$ and such that $C$ is one of its $f$-separating cms’s. If the $f$-component $H$ of $C$ is the graph in Figure 8(a) such that $[f]$ is $w_{1}w_{2}$ and $C$ is the subgraph induced by $\{w_{3},w_{4},w_{5},w_{6}\}$, then we can use Lemma 4 Statement (2) to show that the expanded $f$-component of $C^{-1}$ has the desired properties. Hence, we complete the proof by showing that $H$ has these properties. We prove this by first showing that (i) $H$ has no $[f]$-preserving forbidden minor. Second, we use (i) to show that (ii) some minor $[H]_{A}$ is the graph in Figure 8(a) such that $[f]_{A}$ is $w_{1}w_{2}$ and $|[y]_{A}^{-1}|=1$ for each vertex $y\in V(C)$. Third and finally, letting $f=uv$, we use (i) and (ii) to show that (iii) $|[u]_{A}^{-1}|=|[v]_{A}^{-1}|=1$. This completes the proof.
For Statement (i), since $H$ is a $C$-component of $[G\cup f]$, Lemma 60 in Appendix E states that any atom $J$ of $H$ is also an atom of $[G\cup f]$. Additionally, by Lemma 33, $J$ contains $[f]$. Since $[G\cup f]$ is $f$-separating, these facts imply that $J$ has no $[f]$-preserving forbidden minor. Therefore, Statement (i) follows from the contrapositive of Lemma 27 in Appendix C.2.
For Statement (ii), note that $C$ does not contain any endpoint of $[f]$, or else we can use Lemma 4 Statement (2) to show that $H$ has an $[f]$-preserving $K_{5}$ minor, which contradicts Statement (i). Using the same lemma, we see that $H$ has a $K_{5}$ minor $[H]_{B}$ such that $|[y]_{B}^{-1}|=1$, for each vertex $y\in V(C)$. By Statement (i), $f$ is contracted in $[H]_{B}$. Hence, there exists a minor $[H]_{A}$ from which $[H]_{B}$ can be obtained by contracting $[f]_{A}$. Clearly, $[H]_{A}$ is one of the graphs in Figures 8(a), 17(a), or 18(a)-18(d) such that $[f]_{A}$ is $w_{1}w_{2}$ and $[C]_{A}$ is the subgraph induced by $\{w_{3},w_{4},w_{5},w_{6}\}$. Since $G\cup f$ is an atom, $H$ is $2$-connected. Therefore, Lemma 10 allows us to assume that $[H]_{A}$ is one of the graphs in Figures 8(a) or 18(a)-18(d). Statement (ii) now follows from Statement (i) and Lemma 11.
Finally, for Statement (iii), wlog assume to the contrary that $|[u]_{A}^{-1}|>1$. If either $[u]$ has a neighbor in $C$ or $|[u]_{A}^{-1}|>2$, then we can use Lemma 5 to contradict Statement (i). Otherwise, let $[x]$ be the unique vertex in $[u]_{A}^{-1}\setminus\{[u]\}$. If $|[v]_{A}^{-1}|=1$, then $\{[x],[v]\}\}$ is a clique separator of $H$. However, since neither $[x]$ nor $[v]$ is contained in $C$, we have $1=|[x]^{-1}|=|[v]^{-1}|$. This implies that $G\cup f$ is not as atom, which is a contradiction. Hence, assume that $|[v]_{A}^{-1}|>1$. Similar arguments show that $[v]$ has no neighbor in $C$, there exists a unique vertex $[x^{\prime}]$ in $[v]_{A}^{-1}\setminus\{[v]\}$, and $1=|[x]^{-1}|=|[x^{\prime}]^{-1}|$. Hence, $\{x,x^{\prime}\}$ is a size two separator of $G\cup f$, and it is easy to see that in any $f$-preserving forbidden minor of $G\cup f$ some edge $e$ in the set $\{ux,ux^{\prime},vx,vx^{\prime}\}$ is contracted. Furthermore, the minor obtained by contracting $e$ has an $f$-separating pair whose $f$-component $J$ is such that $J^{-1}$ is a proper subgraph of $H^{-1}$. This contradicts the fact that $C$ is in the top-level of $[G\cup f]$, and so Statement (iii) is proved. This completes the proof. ∎
@@
By Lemma 4 Statement (1), $C^{-1}$ is one of the graphs in Figure 13. We will show that $C^{-1}$ is not any of the graphs in Figures 13(e)-(i). Assume to the contrary that $C^{-1}$ is one of these graphs. We will show that $(G,f)$ has a reducing edge, and hence is not minimal, which is a contradiction. Let $H^{-1}$ and $I^{-1}$ be the expanded $f$-component and expanded minor component of $C^{-1}$, respectively. By Lemma 34, $H^{-1}$ is one of the graphs in Figure 19 such that $f$ is the red edge and $C^{-1}$ is the subgraph induced by the set $\{w_{i}\}$. We consider each case for $H^{-1}$.
@@
If $H^{-1}$ is the graph in either Figure 19(a) or 19(b), then let $[G\cup f]$ be the minor obtained by contracting $uw_{1}$. Observe that $f$ is retained in $[G\cup f]$ and $[H^{-1}]$ is an atom that contains $[f]$ such that $[H^{-1}]\cup[w_{1}w_{4}]$ has an $f$-preserving forbidden minor. Hence, some atom $J$ of $[G\cup f]$ contains $[H^{-1}]$. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. Since $G^{\prime}$ is preserved in $[G\cup f]$, Lemma 4 Statement (2) tells us that some path in $[G^{\prime}]$ between $[w_{1}]$ and $[w_{4}]$ does not contain any vertex in $[C^{-1}]\setminus\{[w_{1}],[w_{4}]\}$. Note that the only vertices in this path that are also in $[H^{-1}]$ are $[w_{1}]$ and $[w_{4}]$. Therefore, Lemma 66 shows that some path in $J$ between $[w_{1}]$ and $[w_{4}]$ does not contain any vertex in $[H^{-1}]\setminus\{[w_{1}],[w_{4}]\}$. The above facts imply that $J$ has an $[f]$-preserving forbidden minor, and so $uw_{1}$ is reducing, as desired.

### v2

@@ B.1   Proof of Proposition 6
By Lemma 19, every graph-nonedge pair in Figure 16 has a proper non-$3$-SIP map. It remains to show that these are the only minimal graph-nonedge pairs with no Type (2) edge. Consider any graph-nonedge pair $(G,f)$, where $f=uv$. If $(G,f)$ is a pair in Figure 16, then it is easy to verify that $(G,f)$ is minimal and has no Type (2) edge. Next, we show that if $(G,f)$ is minimal and has no Type (2) edge, then it is one of the pairs in Figure 16. Since $(G,f)$ is minimal, $G\cup f$ is an atom and has an $f$-preserving forbidden minor $[G\cup f]$. Furthermore, Statements (i)-(iv) of Lemma 20 are true and show that $5\leq|V(G)|\leq 8$. Let $H=[G\cup f]\setminus\{[u],[v]\}$ and consider the subgraph $H^{-1}$ of $G\cup f$. If $|[u]^{-1}|=|[v]^{-1}|=1$, then Statement (ii) shows that $(G,f)$ is the pair in either Figure 16(a) or 16(b). Hence, assume wlog that $|[u]^{-1}|>1$. By Statement (iv), we have $[u]^{-1}=\{u,x\}$. There are three cases.
@@
Let $w$ be the vertex in $H^{-1}$ that neighbors $x$ but not $u$ if such a vertex exists, or any vertex in $H^{-1}$ that neighbors $x$ otherwise. Also, let $J$ be the subgraph of $G\cup f$ induced by $[u]^{-1}$. By Statement (iii), $J\setminus x$ is connected and contains $u$. Hence, $xw$ is contracted in the minor obtained via the vertex exchange from $[u]^{-1}$ to $[w]^{-1}$ that exchanges $x$ and fixes $J\setminus x$, and this minor has an $f$-preserving forbidden minor, contradicting Statement (i).
@@
Next, consider the case where $|[v]^{-1}|=1$. If $xw$ is a nonedge of $G\cup f$, then $(G,f)$ is the pair in either Figure 16(c) or 16(d). Otherwise, $\{x,v,w\}$ is a clique separator of $G\cup f$, contradicting the fact that $G\cup f$ is an atom. Lastly, consider the case where $|[v]^{-1}|>1$. The discussion above shows that $w$ is the only neighbor of $v$ in $H^{-1}$. Let $w^{\prime}$ be any vertex in $H^{-1}\setminus w$ that neighbors $x$. Also, let $J$ be defined as in Case 2. Then, $xw^{\prime}$ is contracted in the minor obtained via the vertex exchange from $[u]^{-1}$ to $[w^{\prime}]^{-1}$ that exchanges $x$ and fixes $J\setminus x$, and this minor has an $f$-preserving forbidden minor, contradicting Statement (i).
@@
B.2   Proofs of lemmas for Proposition 8
We present proofs of the lemmas in Section 5.1 up to Lemma 34, the last one before the proof of Proposition 8. We start with the proof of Lemma 25.
Proof of Lemma 25.
For Statement (1), since $(G\cup f)\setminus C^{-1}$ is disconnected and $G\cup f$ is an atom, $C^{-1}$ is not a clique. Hence, Figure 22 shows all cases for $C^{-1}$ if $|V(C^{-1})|\leq 5$. We complete the proof by showing that $|V(C^{-1})|\leq 5$. Assume to the contrary that this is not the case. Consider any $f$-separating minor $[G\cup f]_{1}$ such that $(C,M)$ is one of its $f$-separating pairs. Then, $C$ is a clique on at least five vertices. Consequently, since any atom $J$ of $[G\cup f]_{1}$ that contains $[f]_{1}$ has no $[f]_{1}$-preserving forbidden minor, by definition, $J$ cannot contain $C$. This implies that there exists a clique minimal $[ux]_{1}$-separator $E$ of $[G\cup f]_{1}$, for some endpoint $[u]_{1}$ of $[f]_{1}$ and any vertex $[x]_{1}$ in $C\setminus E$, such that $C\setminus E$ is non-empty. Clearly, $(E^{-1},M^{-1})$ is an expanded $f$-separating pair of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$. However, this shows that $C^{-1}$ is not in the top-level of $(G,f)$, which is a contradiction. Therefore, we have $|V(C^{-1})|\leq 5$, and so Statement (1) is proved.
Next, for Statement (2), observe that $C^{-1}$ is retained in $[H^{-1}\cup I^{-1}]$, and $[H^{-1}]$ and $[I^{-1}]$ are vertices of $[H^{-1}\cup I^{-1}]$. Hence, Statement (1) shows that $[C^{-1}]$ is one the graphs in Figure 22. Consequently, $[H^{-1}\cup I^{-1}]$ is some graph $G^{\prime}$ that can be obtained from one of the graphs in Figure 23 by deleting some edges, each of which has one endpoint in $\{a,b\}$ and endpoint in $\{w_{i}\}$. Furthermore, $[C^{-1}]$ is the subgraph of $G^{\prime}$ induced by the set $\{w_{i}\}$. If $G^{\prime}$ can be obtained without deleting any edges, then we are done. Otherwise, assume wlog that the above-mentioned minor $[G\cup f]_{1}$ is obtained by contracting $w_{1}w_{2}$. It is easy to see that either $G\cup f$ is not an atom or, for any endpoint $[u]_{1}$ of $[f]_{1}$ and any vertex $[x]_{1}$ of $M\setminus C$, $C$ is not a minimal $[ux]_{1}$-separator of $[G\cup f]_{1}$. In either case, we get a contradiction, and thus Statement (2) is proved. ∎
Lemma 28 is proved using Lemmas 26 and 27 in Section 5.1. The proof of Lemma 26 requires Lemma 42, below.
Let $C^{-1}$ be any expanded $f$-separating CMS of a minimal graph-nonedge pair $(G,f)$, with $f=uv$, and let its expanded $f$-component and expanded minor component be $H^{-1}$ and $I^{-1}$, respectively. Also, consider the minor $[G\cup f]$ obtained by contracting any edge of $(H^{-1}\setminus f)\setminus C^{-1}$ and any $f$-preserving forbidden minor $[I^{-1}]_{1}$. There exists a pair $xy$ of distinct vertices in $C^{-1}$ that is preserved in $[I^{-1}]_{1}$ and some paths $P_{ux}$ and $P_{vy}$ in $G\cup f$ between $u$ and $x$ and between $v$ and $y$, respectively, that each contain exactly one vertex in $I^{-1}$ and such that $[P_{ux}]$ and $[P_{vy}]$ are vertex-disjoint.
@@ B.2   Proofs of lemmas for Proposition 8
Since $f$ is preserved in $[I^{-1}]_{1}$ and $(G\cup f)\setminus C^{-1}$ is disconnected, there exists a pair of distinct vertices $x^{\prime}y^{\prime}$ in $C^{-1}$ and an extension $[G\cup f]_{2}$ of $[I^{-1}]_{1}$ such that $[f]_{2}=[x^{\prime}y^{\prime}]_{2}$. This implies that $x^{\prime}y^{\prime}$ is preserved in $[I^{-1}]_{1}$ and some vertex-disjoint paths $P_{ux^{\prime}}$ and $P_{vy^{\prime}}$ in $G\cup f$ between $u$ and $x^{\prime}$ and between $v$ and $y^{\prime}$, respectively, each contain exactly one vertex in $I^{-1}$. If $[P_{ux^{\prime}}]$ and $[P_{vy^{\prime}}]$ are vertex-disjoint, then setting $xy=x^{\prime}y^{\prime}$ proves the lemma.
@@
By definition, $I^{-1}$ has an $f$-preserving forbidden minor $[I^{-1}]_{1}$. Hence, since $I^{-1}$ is preserved in $[G\cup f]$, it suffices to show that the edge contracted in $[G\cup f]$ is contracted in some extension of $[I^{-1}]_{1}$. This follows easily from Lemma 42. ∎
The proof of Lemma 27 requires Lemma 43, below.
@@
Proof of Lemma 27.
@@
Lemma 30 is proved using Lemma 29 in Section 5.1. The proof of Lemma 29 requires Lemma 44, below.
@@
Consider any $f$-separating minor $[G\cup f]$ obtained by contracting a single edge of $C^{-1}$ and such that $C$ is one of its $f$-separating CMSs. If the $f$-component $H$ of $C$ is the graph in Figure 8(a) with $[f]$ as $w_{1}w_{2}$ and $C$ as the subgraph induced by $\{w_{3},w_{4},w_{5},w_{6}\}$, then we can use Lemma 25 Statement (2) to show that the expanded $f$-component of $C^{-1}$ has the desired properties. Hence, we complete the proof by showing that $H$ has these properties. We prove this by first showing that (i) $H$ has no $[f]$-preserving forbidden minor. Second, we use (i) to show that (ii) some minor $[H]_{1}$ is the graph in Figure 8(a) such that $[f]_{1}$ is $w_{1}w_{2}$ and $|[y]_{1}^{-1}|=1$ for each vertex $y\in V(C)$. Third and finally, letting $f=uv$, we use (i) and (ii) to show that (iii) $|[u]_{1}^{-1}|=|[v]_{1}^{-1}|=1$. This completes the proof.
For Statement (i), since $H$ is a $C$-component of $[G\cup f]$, Lemma 62 in C states that any atom $J$ of $H$ is also an atom of $[G\cup f]$. Additionally, by Lemma 44, $J$ contains $[f]$. Since $[G\cup f]$ is $f$-separating, these facts imply that $J$ has no $[f]$-preserving forbidden minor. Therefore, Statement (i) follows from the contrapositive of Lemma 22.
For Statement (ii), note that $C$ does not contain any endpoint of $[f]$, or else we can use Lemma 25 Statement (2) to show that $H$ has an $[f]$-preserving $K_{5}$ minor, which contradicts Statement (i). Using the same lemma, we see that $H$ has a $K_{5}$ minor $[H]_{2}$ such that $|[y]_{2}^{-1}|=1$, for each vertex $y\in V(C)$. By Statement (i), $f$ is contracted in $[H]_{2}$. Hence, there exists a minor $[H]_{1}$ from which $[H]_{2}$ can be obtained by contracting $[f]_{1}$. Clearly, $[H]_{1}$ is one of the graphs in Figures 8(a), 14(a), 15(a), or 15(b) with $[f]_{1}$ as $w_{1}w_{2}$ and $[C]_{1}$ as the subgraph induced by $\{w_{3},w_{4},w_{5},w_{6}\}$. Since $G\cup f$ is an atom, $H$ is $2$-connected. Therefore, Lemma 4 allows us to assume that $[H]_{1}$ is one of the graphs in Figures 8(a), 15(a), or 15(b). Statement (ii) now follows from Statement (i) and Lemma 5.
Finally, for Statement (iii), wlog assume to the contrary that $|[u]_{1}^{-1}|>1$. If either $[u]$ has a neighbor in $C$ or $|[u]_{1}^{-1}|>2$, then we can use Lemma 28 to contradict Statement (i). Otherwise, let $[x]$ be the unique vertex in $[u]_{1}^{-1}\setminus\{[u]\}$. If $|[v]_{1}^{-1}|=1$, then $\{[x],[v]\}\}$ is a clique separator of $H$. However, since neither $[x]$ nor $[v]$ is contained in $C$, we have $1=|[x]^{-1}|=|[v]^{-1}|$. This implies that $G\cup f$ is not as atom, which is a contradiction. Hence, assume that $|[v]_{1}^{-1}|>1$. Similar arguments show that $[v]$ has no neighbor in $C$, there exists a unique vertex $[x^{\prime}]$ in $[v]_{1}^{-1}\setminus\{[v]\}$, and $1=|[x]^{-1}|=|[x^{\prime}]^{-1}|$. Hence, $\{x,x^{\prime}\}$ is a size two separator of $G\cup f$, and it is easy to see that in any $f$-preserving forbidden minor of $G\cup f$ some edge $e$ in the set $\{ux,ux^{\prime},vx,vx^{\prime}\}$ is contracted. Furthermore, the minor obtained by contracting $e$ has an $f$-separating pair whose $f$-component $J$ is such that $J^{-1}$ is a proper subgraph of $H^{-1}$. This contradicts the fact that $C$ is in the top-level of $[G\cup f]$, and so Statement (iii) is proved. This completes the proof. ∎
Proof of Lemma 30.
@@
By Lemma 25 Statement (1), $C^{-1}$ is one of the graphs in Figure 22. We will show that $C^{-1}$ is not any of the graphs in Figures 22(e)-(i). Assume to the contrary that $C^{-1}$ is one of these graphs. We will show that $(G,f)$ has a reducing edge, and hence is not minimal, which is a contradiction. Let $H^{-1}$ and $I^{-1}$ be the expanded $f$-component and expanded minor component of $C^{-1}$, respectively. By Lemma 29, $H^{-1}$ is one of the graphs in Figure 24 with $f$ as the red edge and $C^{-1}$ as the subgraph induced by the set $\{w_{i}\}$. We consider each case for $H^{-1}$.
@@
If $H^{-1}$ is the graph in either Figure 24(a) or 24(b), then let $[G\cup f]$ be the minor obtained by contracting $uw_{1}$. Observe that $f$ is retained in $[G\cup f]$ and $[H^{-1}]$ is an atom that contains $[f]$ such that $[H^{-1}]\cup[w_{1}w_{4}]$ has an $f$-preserving forbidden minor. Hence, some atom $J$ of $[G\cup f]$ contains $[H^{-1}]$. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. Since $G^{\prime}$ is preserved in $[G\cup f]$, Lemma 25 Statement (2) tells us that some path in $[G^{\prime}]$ between $[w_{1}]$ and $[w_{4}]$ does not contain any vertex in $[C^{-1}]\setminus\{[w_{1}],[w_{4}]\}$. Note that the only vertices in this path that are also in $[H^{-1}]$ are $[w_{1}]$ and $[w_{4}]$. Therefore, Lemma 68 in C shows that some path in $J$ between $[w_{1}]$ and $[w_{4}]$ does not contain any vertex in $[H^{-1}]\setminus\{[w_{1}],[w_{4}]\}$. The above facts imply that $J$ has an $[f]$-preserving forbidden minor, and so $uw_{1}$ is reducing, as desired.